OpenStudy (anonymous):

Solve equation log(2x + 1) = 1 + log (x-2)

OpenStudy (anonymous):

The book said 21/8

OpenStudy (xapproachesinfinity):

so do you know any log properties?

OpenStudy (xapproachesinfinity):

\[\large \log a-\log b=?\]

OpenStudy (anonymous):

a/b

OpenStudy (xapproachesinfinity):

you mean \[=\log[\frac{a}{b}]\] yes?

OpenStudy (dumbcow):

you may need to specify which base the log is in, natural log or base 10 ?

OpenStudy (xapproachesinfinity):

and yes we need to know what base is it!!!

OpenStudy (anonymous):

yes, I used that

OpenStudy (xapproachesinfinity):

so are we dealing with natural log or base 10?

OpenStudy (dumbcow):

from the answer given, looks like base 10

OpenStudy (xapproachesinfinity):

right it has to be log base 10

OpenStudy (xapproachesinfinity):

we need to subtract log(x-2) from both sides to have two logs in one side so we can use our property above! okay can you do that

OpenStudy (xapproachesinfinity):

do it on the paper and show me what you got!!

OpenStudy (anonymous):

(2x+1)/(x-2)=10

OpenStudy (xapproachesinfinity):

show all your work!

OpenStudy (xapproachesinfinity):

how did you get that?! it's right but show your work

OpenStudy (xapproachesinfinity):

oh well just finish solving that equation...

OpenStudy (anonymous):

still